\begin{thebibliography}{1}
\bibitem{test} Aczél, J., Dhombres, J., \textit{Functional equations in several variables with applications to mathematics, Information theory and to the natural and social sciences. Encyclopedie of Mathematics and its Applications, vol. 31, Cambridge University Press, Cambridge (34B40) MR1004465 (90h: 39001) (1989).}
\bibitem{Ajb} Ajebbar, O., Elqorachi, E., \textit{D'Alembert's other functional equation on semigroups.
Aequat.Math, \textbf{94} 913--930 (2020).}
\bibitem{BE} Bouikhalene, B., Elqorachi, E., \textit{A class of functional equations of type d'Alembert on monoids. In : Anastassiou, G., Rassias, J.M. (eds.) Frontiers in Functional Equations and Analytic Inequalities \textbf{pp} 219--235. Springer, Berlin (2019).}
\bibitem{D1}  D'Alembert, J., \textit{Recherche sur la courbe que forme une corde tendue mise en vibration.
History of the Academy, vol I, 1747, Berlin \textbf{pp} 214--219 (1747).}
\bibitem{D2}  D'Alembert, J., \textit{Recherche sur la courbe que forme une corde tendue mise en vibration.
History of the Academy, vol II, 1747, Berlin \textbf{pp} 220--249 (1747).}
\bibitem{ES1}  Ebanks, B., \textit{The sine addition and subtraction formulas on semigroups.
Acta Math. Hungar, \textbf{164}(2) 533--555 (2021).}
\bibitem{EB3}  Ebanks, B., \textit{D'Alembert's other functional equation on monoids revisited.
Publ. Math. Debrecen, (2021).}
\bibitem{EB2}  Ebanks, B., \textit{Around the sine addition law and
d’Alembert’s equation on semigroups.
Results Math, \textbf{77}, 11 (2022). https://doi.org/10.1007/s00025-021-01548-6}
\bibitem{ES} Ebanks, B., Stetkaer, H., \textit{D'Alembert's other functional equation on monoids with involution. Aequat.Math, \textbf{89} 187--206 (2015).}
\bibitem{EBA} Ebanks, B., \textit{The Cosine-Sine functional equation on semigroups. Ann. Math, vol.36, no.1, 2022, \textbf{pp} 30-52. https://doi.org/10.2478/amsil-2021-0012}
\bibitem{K} Kannappan, Pl., \textit{Functional equations and inequalities with applications. Springer monographs in Mathematics, Springer, New York, (2009) https://doi.org/10.1007/987-0-387-89492-8}
\bibitem{ST} Stetkaer, H., \textit{Functional equations on Groups.
World scientific Publishing CO, Singapore (2013).}
\end{thebibliography}

\documentclass{birkjour}
 \newtheorem{thm}{Theorem}[section]
 
 \newtheorem{lem}[thm]{Lemma}
 \newtheorem{prop}[thm]{Proposition}
 \theoremstyle{definition}
 \newtheorem{defn}[thm]{Definition}
 \theoremstyle{remark}
 \newtheorem{rem}[thm]{Remark}
 
 \numberwithin{equation}{section}

\begin{document}

%-------------------------------------------------------------------------
% editorial commands: to be inserted by the editorial office
%
%\firstpage{1} \volume{228} \Copyrightyear{2004} \DOI{003-0001}
%
%
%\seriesextra{Just an add-on}
%\seriesextraline{This is the Concrete Title of this Book\br H.E. R and S.T.C. W, Eds.}
%
% for journals:
%
%\firstpage{1}
%\issuenumber{1}
%\Volumeandyear{1 (2004)}
%\Copyrightyear{2004}
%\DOI{003-xxxx-y}
%\Signet
%\commby{inhouse}
%\submitted{March 14, 2003}
%\received{March 16, 2000}
%\revised{June 1, 2000}
%\accepted{July 22, 2000}
%
%
%
%---------------------------------------------------------------------------
%Insert here the title, affiliations and abstract:
%

%----------Author 1
\author[Youssef Aserrar and Elhoucien Elqorachi]{Youssef Aserrar and Elhoucien Elqorachi}

\address{% 
	Ibn Zohr University, Faculty of sciences, 
Department of mathematics,\\
Agadir,
Morocco}

\email{youssefaserrar05@gmail.com, elqorachi@hotmail.com }
%----------classification, keywords, date
\subjclass{39B52, 39B32}

\keywords{Semigroups, Prime ideal, Involutive automorphism, Multiplicative function, d'Alembert's equation.}

\date{January 1, 2004}
%----------additions
%\dedicatory{To my boss}
%%% ----------------------------------------------------------------------
\title[A d'Alembert type functional equation on  semigroups]{A d'Alembert type functional equation on semigroups}
 
\begin{abstract}
We treat two related trigonometric functional equations on
semigroups. First we solve the $\mu$-sine subtraction law
\[\mu(y) k(x \sigma(y))=k(x) l(y)-k(y) l(x), \quad x, y \in S,\]
for $k, l : S\rightarrow \mathbb{C}$, where $S$ is a semigroup  and 
 $\sigma$ an involutive automorphism, $\mu :S\rightarrow \mathbb{C}$ is a multiplicative function  such that
$\mu (x\sigma (x))=1$ for all $x\in S$, then we determine the complex-valued solutions of the
following functional equation 
\[f(xy) - \mu (y)f(\sigma (y)x) = g(x)h(y),\quad x,y\in S,\] 
on a larger class of semigroups.
\end{abstract}

%%% ----------------------------------------------------------------------
\maketitle

%%% ----------------------------------------------------------------------
%\tableofcontents
\section{Introduction}
Ebanks and Stetkaer \cite{ES} solved the functional equation
\begin{equation}
f(xy) -f(\sigma (y)x) = g(x)h(y),\quad x,y\in M,
\label{EB1}
\end{equation} 
for $f,g,h: M \rightarrow \mathbb{C}$, where $M$ is a group or a monoid generated by its squares and $\sigma :M\rightarrow M$ is an involutive automorphism. That is $\sigma (xy)=\sigma (x)\sigma (y)$ and $\sigma(\sigma(x))=x$ for all $x,y\in M$.\newline
In \cite{BE},  Bouikhalene and Elqorachi determined the complex-valued solutions of the functional equation
\begin{equation}
f(x y)-\mu(y) f(\sigma(y) x)=g(x) h(y), \quad x, y \in M,
\label{E1}
\end{equation}
on groups and monoid generated by its squares, where $\mu :M\rightarrow \mathbb{C}$ is a multiplicative function such that $\mu (x\sigma (x))=1$ for all $x\in M$.
Moreover, the solutions of \eqref{E1} on a semigroup generated by its squares are also known (See Ajebbar and Elqorachi \cite{Ajb}). Recently,  Ebanks \cite{EB3} solved \eqref{EB1} on a larger class of monoids that contains the class of monoids generated by their squares, and regular monoids.\\
 There are some results about solutions of Equation \eqref{EB1} on abelian groups in the literature. The case of $G=\mathbb{R}$ is treated by Kannappan \cite[Equation (vs) in section 3.4.9]{K}. Stetkaer \cite[Corollary 3.5]{ST} derives the solution formulas for the functional equation \eqref{EB1} on abelian groups. The method used in \cite{Ajb,BE,EB3,ES} proceeds from the starting point of the trigonometric functional equation
\begin{equation}
\mu(y) k(x \sigma(y))=k(x) l(y)-k(y) l(x), \quad x, y \in S.
\label{E2}
\end{equation}
The solutions of \eqref{E2} (with $\mu=1$) on a general monoid was given recently by Ebanks \cite[Theorem 4.2]{ES1} in terms of multiplicative and additive functions. In the present paper, we treat the functional equation \eqref{E1} on a general class of semigroups.\\
The contributions of the present work to the knowledge about solutions of \eqref{E1} are the following :
\begin{enumerate}
\item[(1)] The setting has $S$ to be a semigroup, not necessarily generated by its squares, or a monoid.
\item[(2)] We solve Equation \eqref{E2} on semigroups.
\item[(3)] We relate the solutions of \eqref{E1} to the functional equation \eqref{E2} on semigroups, and find explicit formulas for the solutions on a larger class of semigroups.
\end{enumerate}

The outline of the paper is as follows : In the next section we give notations ans terminology. In the third section we solve the functional equation \eqref{E2} on semigroups, and we obtain the explicit formulas for the solutions of Equation \eqref{E1} on a larger class of semigroups $S$.

\section{Notations and Terminology}
Throughout this paper $S$ denotes a semigroup.\\
A function $A: S\rightarrow \mathbb{C}$ is additive if $A(xy) = A(x) + A(y)$ for all $x, y\in S$.\\
A function $\chi: S\rightarrow \mathbb{C}$ is multiplicative if $\chi(xy) = \chi(x)\chi(y)$ for all $x, y \in S$.\\
A function $f: S\rightarrow \mathbb{C}$ is central if $f(xy) = f(yx)$ for all $x, y\in S$, and $f$ is abelian if $f$ is central and $f(xyz)=f(xzy)$ for all $x,y,z\in S.$\\
If $S$ is a semigroup, we define the nullspace
$$
\mathcal{N}_{\mu}(\sigma, S):=\left\lbrace \theta: S \rightarrow \mathbb{C} \mid \theta(x y)-\mu(y) \theta(\sigma(y) x)=0,\quad x, y \in S\right\rbrace .
$$
For any subset $T\subseteq S$ we define $T^2:=\lbrace xy\  \vert \ x, y\in T\rbrace$.
If $\chi: S \rightarrow \mathbb{C}$ is a multiplicative function and $\chi \neq 0$, we define the sets $$I_{\chi}:=\{x \in S \mid \chi(x)=0\}$$ 
$$P_\chi : =\{p\in I_{\chi}\backslash I_{\chi}^2\  \vert up, pv, upv\in I_{\chi}\backslash I_{\chi}^2\  \text{for all}\  u,v\in S\backslash I_{\chi}\}.$$
 Notice that $I_{\chi}$ is  a subsemigroup of $S$ and a prime ideal of $S$.\par  
 For any function $f: S \rightarrow \mathbb{C}$ we define the function $f^{*}(x)=\mu(x) f(\sigma(x))$,\\ $x \in S$, we call $f^{e}:=\frac{f+f^{*}}{2}$ the even part of $f$ and $f^{\circ}:=\frac{f-f^{*}}{2}$ its odd part. The function $f$ is said to be even if $f=f^{*}$, and $f$ is said to be odd if $f=-f^{*}$. If $g, h: S \rightarrow \mathbb{C}$ are two functions we define the function $$(g \otimes h)(x, y):=g(x) h(y),\ x, y \in S.$$
  For a topological semigroup $S$ let $C(S)$ denote the algebra of continuous functions from $S$ into $\mathbb{C}$.
\section{Main result}
The following Lemma will be used later.
\begin{lem}
Let $k, l : S\rightarrow\mathbb{C}$ be a solution of the functional equation \eqref{E2} with $k \neq 0$. Then 
\begin{enumerate}
\item[(1)] $k(xy)=-k^*(yx)$ for all $x,y\in S$.
\item[(2)] If $k=k^*$, then $k(xyz)=0$ for all $x,y,z\in S$.
\end{enumerate}
\label{lem}
\end{lem}
\begin{proof}
(1) Let $x,y \in S$ be arbitrary. By interchanging $x$ and $y$ in \eqref{E2} we get that $\mu (y) k\left( x\sigma (y)\right) =-\mu (x) k\left( y\sigma (x)\right)$, and if we apply this identity to the pair $\left(x, \sigma (y) \right) $, we get $\mu \left( \sigma (y)\right)  k\left( xy\right)=-\mu (x) k\left( \sigma (yx)\right)$.\\
Multiplying this by $\mu (y)$ and by using that $\mu : S\rightarrow \mathbb{C}$ is a multiplicative function such that $\mu \left(y\sigma (y) \right) =1$ for all $y\in S$, we get $k(xy)=-k^*(yx)$ for all $x,y \in S$.\\
(2) By using (1), we get 
$$k(xyz)=-k^*(zxy)=k(yzx)=-k^*(xyz),\quad \text{for all}\quad x,y,z\in S.$$
So if $k=k^*$, we get that $k(xyz)=-k(xyz)$ for all $x,y,z\in S$. This implies that $k(xyz)=0$. This completes the proof of Lemma \ref{lem}.
\end{proof}
In the following proposition we extend the results obtained by Ajebbar and Elqorachi \cite[Proposition 3.3]{Ajb} on  semigroups generated by their squares and by Ebanks \cite[Theorem 4.2]{ES1} on monoids to general semigroups.
\begin{prop}
 The solutions of the functional equation \eqref{E2} with $k \neq 0$ are the following pairs :
 \begin{enumerate}
 \item[(1)] $k$ is any non-zero function such that $k=0$ on $S^2$ and $l=ck$, where $c \in \mathbb{C}$ is a constant.
\item[(2)] $$k=c_1\dfrac{\chi -\chi ^*}{2}\quad\text{and}\quad l=\dfrac{\chi +\chi ^*}{2}+c_2\dfrac{\chi -\chi ^*}{2},$$
where $\chi  : S\rightarrow \mathbb{C}$ is a multiplicative function such that $\chi ^*\neq \chi$ and $c_1 \in \mathbb{C}\backslash  \lbrace0\rbrace,c_2\in \mathbb{C}$ are constants.
\item[(3)] $$k=\left\{ \begin{matrix}
   \chi A & on & S\backslash {{I}_{\chi }}  \\
   0 & on & {{I}_{\chi }}\backslash {{P}_{\chi }}  \\
   \rho  & on & {{P}_{\chi }}  \\
\end{matrix} \right.\quad\text{and}\quad l=\left\{ \begin{matrix}
   \chi (1+cA) & on & S\backslash {{I}_{\chi }}  \\
   0 & on & {{I}_{\chi }}\backslash {{P}_{\chi }}  \\
   c\rho  & on & {{P}_{\chi }}  \\
\end{matrix} \ ,\right.
$$ 
\end{enumerate}
where $c \in \mathbb{C}$ is a constant, $\chi: S \rightarrow \mathbb{C}$ is a non-zero multiplicative function and $A: S \backslash I_{\chi} \rightarrow \mathbb{C}$ is a non-zero additive function such that $\chi^{*}=\chi$, $A \circ \sigma=-A$, and $\rho: P_{\chi} \rightarrow \mathbb{C}$ is the restriction of $k$ to $P_\chi$ such that $\rho^*=-\rho$. In addition we have the following conditions: \\
(I): If $x\in\left\lbrace up,pv,upv \right\rbrace $ for $p\in P_{\chi}$ and $u,v\in S\backslash I_{\chi}$, then $x\in P_{\chi}$ and we have respectively $\rho(x)=\rho(p)\chi (u)$, $\rho(x)=\rho(p)\chi (v)$, or $\rho(x)=\rho(p)\chi (uv)$.\\
(II): $k(xy)=k(yx)=0$ for all $x\in S\backslash I_{\chi}$ and $y\in I_{\chi}\backslash P_{\chi}$.\par
Conversely, if $(k,l)$ are given by the formulas in (1), (2) or (3) with conditions (I) and (II) holding and if $k$ is abelian in case (3), then $(k,l)$ satisfies \eqref{E2}.\par
 Furthermore, if $S$ is a topological semigroup and $k\in C(S)$, then \\$\chi,\chi^*\in C(S)$, $A\in C(S\backslash I_{\chi})$, and $\rho \in C(P_{\chi})$.
\label{P1}
\end{prop}
\begin{proof}
We check by elementary computations that if $k$ and $l$ are of the forms (1)--(3) then $(k,l)$ is a solution of \eqref{E2}.\par
Conversely, we will discuss two cases according to whether $k$ and  $l$ are linearly dependent or not.\newline
\underline{First case :} $k$ and $l$ are linearly dependent. There exists $c\in \mathbb{C}$ such that $l=ck$. From \eqref{E2} we get 
\begin{equation}
\mu (y)k(x\sigma (y))=ck(x)k(y)-ck(y)k(x)=0,\quad\text{for all}\quad x,y \in S.
\label{E7}
\end{equation}
 Since $\mu(y\sigma(y))=1$ for all $y\in S$, we deduce from \eqref{E7} that $k=0$ on $S^2$. This occurs in (1) of Proposition \ref{P1}.\newline
\underline{Second case :} $k$ and $l$ are linearly independent. Since $\mu (y\sigma(y))=1$ for all $y\in S$, we multiply \eqref{E2} by $\mu(\sigma(y))$ and we get 
\begin{equation}
k(x\sigma(y))=\mu(\sigma(y))k(x)l(y)-\mu(\sigma(y))k(y)l(x),\quad \text{for all}\quad x,y\in S.
\label{D1}
\end{equation}
By using the associativity of the semigroup operation, we can compute $k(x\sigma(y)\sigma(z))$ first as $k((x\sigma(y))\sigma(z))$ and then as $k(x(\sigma(y)\sigma(z)))$ under the identity \eqref{D1} and compare the results to obtain 
\begin{multline}
-k(z)l(x\sigma(y))=k(x)\left[\mu(\sigma(y))l(yz)-\mu(\sigma(y))l(y)l(z) \right] +\\l(x)\left[\mu(\sigma(y))k(y)l(z)-\mu(\sigma(y))k(yz) \right]. 
\label{D2}
\end{multline}
Since $k\neq 0$, there exists $z_0\in S$ such that $k(z_0)\neq 0$, and then we deduce from the identity above that 
\begin{equation}
l(x\sigma(y))=k(x)f(y)+l(x)g(y),
\label{D3}
\end{equation}
where \\
$$f(y)=\dfrac{-1}{k(z_0)}\left[\mu(\sigma(y))l(yz_0)-\mu(\sigma(y))l(y)l(z_0) \right], $$
and 
\begin{equation}
g(y)=\dfrac{1}{k(z_0)}\left[\mu(\sigma(y))k(yz_0)-\mu(\sigma(y))k(y)l(z_0) \right]. 
\label{D4}
\end{equation}
By substituting \eqref{D3} into \eqref{D2}, we get
\begin{align*}
k(x)(-k(z)f(y))+l(x)(-k(z)g(y))=k(x)\left[\mu(\sigma(y))l(yz)- \mu(\sigma(y))l(y)l(z)\right]+\\l(x)\left[\mu(\sigma(y))k(y)l(z)-\mu(\sigma(y))k(yz) \right].  
\end{align*}
Using the linear independence of $k$ and $l$, we obtain 
$$\mu(\sigma(y))l(yz)=\mu(\sigma(y))l(y)l(z)-k(z)f(y),$$
and 
\begin{equation}
\mu(\sigma(y))k(yz)=\mu(\sigma(y))k(y)l(z)+k(z)g(y).
\label{D5}
\end{equation}
From \eqref{E2} and \eqref{D4} we deduce that the function $g(y)$ can be written as 
$$g(y)=\alpha \mu(\sigma(y))k(y)+\beta \mu(\sigma(y))l(y),$$
where 
$$\alpha =\dfrac{\mu(\sigma(z_0))l(\sigma(z_0))-l(z_0)}{k(z_0)}\quad \text{and}\quad \beta =\dfrac{-\mu(\sigma(z_0))k(\sigma(z_0))}{k(z_0)}.$$
From the last form of $g(y)$ and taking into account that $\mu (y\sigma(y))=1$ for all $y\in S$, equation \eqref{D5} can be written as follows
\begin{equation}
k(yz)=k(y)\left[l(z)+\alpha k(z) \right]+\beta k(z)l(y).
\label{Z1}
\end{equation} 
On the other hand from equation \eqref{D1} we get 
\begin{equation}
k(yz)=k(y)l^*(z)-k^*(z)l(y),
\label{D7}
\end{equation}
then by comparing \eqref{Z1} and \eqref{D7} and using the linear independence of $k$ and $l$ we obtain 
\begin{equation}
l(z)+\alpha k(z)=l^*(z),
\end{equation}
\begin{equation}
\beta k(z)=-k^*(z),\quad \text{for all}\quad z\in S.
\label{D6}
\end{equation}
Since $k\neq 0$, then we get from \eqref{D6} that $\beta \neq 0$ and $\beta^2 =1$. This means that $k=k^*$ or $k=-k^*$. If $k=k^*$ then according to Lemma \ref{lem} (2), $k(xyz)=0$ for all $x,y,z\in S$. This implies that \begin{equation}
k(x)l(yz)=k(yz)l(x),\quad \text{for all}\quad x,y,z\in S.
\label{err1}
\end{equation}  
Since $k$ and $l$ are linearly independent, then $k\neq 0$ on $S^2$, so there exists $y_0,z_0\in S$ such that $k(y_0z_0)\neq 0$. By letting $y=y_0$ and $z=z_0$ in \eqref{err1}, we obtain $l=b k$ for some constant $b\in \mathbb{C}$. This contradicts the fact that $k$ and $l$ are linearly independent. Now if $k=-k^*$ then by using the same computations as in the proof of \cite[Theorem 2.1]{BE}, we get $l^{\circ}=ck$  where $c\in \mathbb{C}$ is a constant and that the pair $(k,l^e)$ satisfies the sine addition law
\[k(xy)=k(x)l^e(y)+k(y)l^e(x),\quad x,y\in S.\]
Hence according to \cite[Theorem 3.1]{EB2} and taking into account that $k\neq 0$, the pair falls into two categories:\\
(i) $k=c_1\dfrac{\chi _1-\chi _2}{2}$ and $l^e=\dfrac{\chi _1+\chi _2}{2}$ , where $\chi _1, \chi _2 : S\rightarrow \mathbb{C}$ are different multiplicative functions and $c_1 \in \mathbb{C}\backslash \lbrace0\rbrace$ is a constant. Since $l=l^e+l^{\circ}$ and $k=-k^*$ we deduce that 
\[k=c_1\dfrac{\chi -\chi ^*}{2}\quad\text{and}\quad l=\dfrac{\chi +\chi ^*}{2}+c_2\dfrac{\chi -\chi ^*}{2},\]
where $\chi  : S\rightarrow \mathbb{C}$ is a multiplicative function such that $\chi ^*\neq \chi$ and $c_1 \in \mathbb{C}\backslash  \lbrace0\rbrace,c_2\in \mathbb{C}$ are constants. This occurs in part (2) of Proposition \ref{P1}.\\
(ii)$$k=\left\{ \begin{matrix}
   \chi A & on & S\backslash {{I}_{\chi }}  \\
   0 & on & {{I}_{\chi }}\backslash {{P}_{\chi }}  \\
   \rho  & on & {{P}_{\chi }}  \\
\end{matrix} \right.\quad\text{and}\quad l^e=\chi \ ,
$$ 
where $\chi: S \rightarrow \mathbb{C}$ is a non-zero multiplicative function and $A: S \backslash I_{\chi} \rightarrow \mathbb{C}$ is a non-zero additive function such that $\chi^{*}=\chi$, $A \circ \sigma=-A$, and $\rho: P_{\chi} \rightarrow \mathbb{C}$ is the restriction of $k$ to $P_\chi$ such that $\rho^*=-\rho$ and we have the conditions (I) and (II). Since $l=l^e+l^{\circ}$ and $l^{\circ}=ck$, we can see that
$$ l=\left\{ \begin{matrix}
   \chi (1+cA) & on & S\backslash {{I}_{\chi }}  \\
   0 & on & {{I}_{\chi }}\backslash {{P}_{\chi }}  \\
   c\rho  & on & {{P}_{\chi }}  \\
\end{matrix} \ .\right.
$$ 
This is part (3).\par 
The topological statements are easy to verify. This completes the proof of Proposition \ref{P1}.
\end{proof}
 Now we solve the functional equation \eqref{E1} on a t-compatible semigroup. A notion which was recently introduced by Ebanks \cite[Definition 4.2]{EBA}. 
 \begin{defn}
Let $S$ be a semigroup. We will say that $S$ is compatible, if $S^2=S$ and for every prime ideal $I\subset S$ the following condition holds.
\begin{equation}
\text{For each}\quad q\in I\quad\text{there exists}\quad w_q\in S\backslash I\quad\text{such that}\quad qw_q\in I^2.
\label{Z2}
\end{equation}
We say that a topological semigroup $S$ is t-compatible, if $S=S^2$ and condition \eqref{Z2} holds for every prime ideal $I$.
\end{defn}
The following result shows that the solutions of \eqref{E1} on a t-compatible semigroup  have the same forms found in \cite[Theorem 4.3]{Ajb} for  semigroups generated by their squares.
\begin{thm}
Let $S$ be a t-compatible semigroup, and suppose that  $f,g,h \in C(S)$ satisfy \eqref{E1}. Then $f,g,h$ belong to one of the four families below. In addition $\chi,\chi^*, \theta\in C(S)$, and $A\in C(S\backslash I_{\chi})$.
\begin{enumerate}
\item[(1)] $f=\theta$, $g=0$ and $h$ is arbitrary, where $\theta \in \mathcal{N}_{\mu}(\sigma, S)$.
\item[(2)] $f=\theta$, $g$ is arbitrary and $h=0$, where $\theta \in \mathcal{N}_{\mu}(\sigma, S)$.
\item[(3)] $f=\theta+\alpha \dfrac{\chi +\chi ^*}{2}+\beta \dfrac{\chi -\chi ^*}{2}$ and \newline $g\otimes h=2\left(\beta \dfrac{\chi +\chi ^*}{2}+\alpha \dfrac{\chi -\chi ^*}{2} \right)\otimes \dfrac{\chi -\chi ^*}{2} $, where $\alpha ,\beta \in \mathbb{C}$ are constants, $\theta \in\mathcal{N}_{\mu}(\sigma, S)$ and $\chi : S\rightarrow \mathbb{C}$ is a multiplicative function such that $\left( \alpha,\beta\right) \neq (0,0)$ and $\chi \neq \chi ^*$.
\item[(4)] $$
\left\{\begin{array}{l}
f=\theta +\dfrac{\alpha}{2}\chi A +\dfrac{\beta}{4}\chi A^2, \quad g=\alpha \chi+\beta \chi  A ,\quad h=\chi A,\quad \text { on } \quad S \backslash I_{\chi} \\
f=\theta, \quad g=0,\quad h=0 \quad \text { on } \quad I_\chi,
\end{array}\right.
$$
\end{enumerate} 
where $\alpha ,\beta \in \mathbb{C}$ are constants, $\theta \in\mathcal{N}_{\mu}(\sigma, S)$ and $\chi : S\rightarrow \mathbb{C}$ is a non-zero multiplicative function such that $\left( \alpha,\beta\right) \neq (0,0)$, $\chi = \chi ^*$ and $A:S\backslash I_{\chi} \rightarrow \mathbb{C}$ is a non-zero additive function such that $A\circ \sigma =-A$.

\label{T1}
\end{thm}
\begin{proof}
We check by elementary computations that if $f$, $g$ and $h$ are of the forms (1)--(4) then $(f,g,h)$ is a solution of \eqref{E1}.\newline
Let $f,g,h :S\rightarrow \mathbb{C}$ satisfy the functional equation \eqref{E1}. If $g=0$ then $h$ is arbitrary and $f=\theta$ where $\theta \in \mathcal{N}_{\mu}(\sigma, S)$. If $h=0$ then $g$ will be arbitrary and $f=\theta$ where $\theta \in \mathcal{N}_{\mu}(\sigma, S)$. So from now on we assume that $g\neq 0$ and $h\neq 0$. According to  \cite[Lemma 4.2 (1)]{Ajb}, there exists a function $l:S\rightarrow \mathbb{C}$ such that 
$$h(xy)=g^*(x)l(y)-g(y)l^*(x),\quad \text{for all}\quad x,y\in S.$$
This implies that $h(xy)=-h^*(yx)$, so $$h(xyz)=-h^*(zxy)=h(yzx)=-h^*(xyz),$$ for all $x,y,z \in S$. Since $S^2=S$ we deduce that $h=-h^*$ and then accordding to \cite[Lemma 4.2 (6)]{Ajb}, there exists a constant $b\in \mathbb{C}$ such that $g^{\circ}=bh$, so we discuss the two cases : $b=0$ and $b\neq 0$.\newline
\underline{First case :} $b\neq 0$. According to the proof of \cite[Theorem 4.3 case A]{Ajb}  there exists a function $m:S\rightarrow \mathbb{C}$ and a constant $c_3\in \mathbb{C}$ such that the pair $(h,m)$ satisfies the $\mu$-sine subtraction law \eqref{E2}, and $g^e=c_3m^e$, then according to Proposition \ref{P1} and taking into account that $h$ is a non-zero odd function, the pair $(h,m)$ falls into the categories:\\
(i) $h=c_1\dfrac{\chi -\chi ^*}{2}$ and $m=\dfrac{\chi +\chi ^*}{2}+\alpha \dfrac{\chi -\chi ^*}{2}$ where $c_1 \in \mathbb{C}\backslash \lbrace 0\rbrace, \alpha \in \mathbb{C}$ are constants and $\chi : S\rightarrow \mathbb{C}$ is a multiplicative function such that $\chi \neq \chi ^*$, then by using similar computations to the ones of the proof of \cite[Theorem 4.3 case A (i)]{Ajb}, we get that 
\[g=c_3\dfrac{\chi +\chi ^*}{2}+c_2\dfrac{\chi -\chi ^*}{2}\quad \text{and}\quad f=\theta +\dfrac{c_1}{2}\left(c_2\dfrac{\chi +\chi ^*}{2}+c_3\dfrac{\chi -\chi ^*}{2} \right), \]
where $c_2=bc_1 \in \mathbb{C}$ is a constant and $\theta \in \mathcal{N}_{\mu}(\sigma, S)$. This occurs in part (3) with $\alpha=\dfrac{c_1c_2}{2}$ and $\beta =\dfrac{c_1c_3}{2}$.\\
(ii) $$h=\left\{ \begin{matrix}
   \chi A & on & S\backslash {{I}_{\chi }}  \\
   0 & on & {{I}_{\chi }}\backslash {{P}_{\chi }}  \\
   \rho  & on & {{P}_{\chi }}  \\
\end{matrix} \right.\quad\text{and}\quad m=\left\{ \begin{matrix}
   \chi (1+cA) & on & S\backslash {{I}_{\chi }}  \\
   0 & on & {{I}_{\chi }}\backslash {{P}_{\chi }}  \\
   c\rho  & on & {{P}_{\chi }}  \\
\end{matrix} \ ,\right.
$$ 
where $c \in \mathbb{C}$ is a constant, $\chi: S \rightarrow \mathbb{C}$ is a non-zero multiplicative function and $A: S \backslash I_{\chi} \rightarrow \mathbb{C}$ is a non-zero additive function such that $\chi^{*}=\chi$, $A \circ \sigma=-A$,  and $\rho: P_{\chi} \rightarrow \mathbb{C}$ is a function such that $\rho^*=-\rho$ and $\rho$ satisfy the condition (I) from Proposition \ref{P1}.\newline
Let $y\in I_{\chi}$. Since $I_{\chi}$ is a prime ideal (if nonempty) and $S$ is t-compatible, then either $y\in I_{\chi}^2$ or there exists $w_y \in S\backslash I_{\chi}$ such that $yw_y \in I_{\chi}^2$. It follows  from condition (I) that $h(y)=0$. Proceeding exactly as in the proof of \cite[Theorem 4.3 case A (ii)]{Ajb}, we find that
\[g=c_3\chi +c_2 \chi A\quad\text{on}\quad S \backslash I_{\chi}\quad\text{and}\quad g=0\quad\text{on}\quad I_\chi,\]
and 
\[f=\theta +\dfrac{c_3}{2}\chi A+\dfrac{c_2}{4} \chi A^2\quad\text{on}\quad S \backslash I_{\chi}\quad\text{and}\quad f=\theta\quad\text{on}\quad I_\chi,\]
where $c, c_{2},c_3 \in \mathbb{C}$ are constants, $\chi: S \rightarrow \mathbb{C}$ is a non-zero multiplicative function and $A: S \backslash I_{\chi} \rightarrow \mathbb{C}$ is a non-zero additive function such that $\chi^{*}=\chi$, $A \circ \sigma=-A$ and $\theta \in \mathcal{N}_{\mu}(\sigma, S)$. This is part (4) of Theorem \ref{T1} with $\alpha=c_3$ and $\beta =c_2$.\\
\underline{Second case :} $b=0$. According to the proof of \cite[Theorem 4.3 case B]{Ajb} there exists  a constant $\lambda \in \mathbb{C}\backslash \lbrace 0\rbrace$ such that the pair $\left( h,\dfrac{\lambda}{2}g\right) $ satisfies the sine addition law  
\begin{equation}
h(xy)=h(x)\left( \dfrac{\lambda}{2}g(y)\right) +h(y)\left( \dfrac{\lambda}{2}g(x)\right), 
\end{equation}
for all $x,y \in S$. Then according to \cite[Theorem 3.1]{EB2} and taking into account that $h$ is a non-zero odd function  and $g$ is even, we have the following possibilities:\\
(i) $h=c_1\dfrac{\chi -\chi ^*}{2}$ and $g=c_2\dfrac{\chi +\chi ^*}{2}$, where $c_1,c_2 \in \mathbb{C}\backslash \lbrace 0\rbrace$ are constants and $\chi : S\rightarrow \mathbb{C}$ is a multiplicative function such that $\chi \neq \chi ^*$, then we get by  using the same computations as in the proof of  \cite[Theorem 4.3 case B (i)]{Ajb} that 
\[f=\theta +\dfrac{c_1c_2}{2}\dfrac{\chi -\chi ^*}{2},\]
where $\theta \in \mathcal{N}_{\mu}(\sigma, S)$. This is part (3) with $\alpha=0$ and $\beta =\dfrac{c_1c_2}{2}$.\\
(ii) $$h=\left\{ \begin{matrix}
   \chi A & on & S\backslash {{I}_{\chi }}  \\
   0 & on & {{I}_{\chi }}\backslash {{P}_{\chi }}  \\
   \rho  & on & {{P}_{\chi }}  \\
\end{matrix} \right.\quad\text{and}\quad g=c_1\chi \  ,
$$ 
where $c_1 \in \mathbb{C}\backslash \left\lbrace 0 \right\rbrace $ is a constant, $\chi: S \rightarrow \mathbb{C}$ is a non-zero multiplicative function and $A: S \backslash I_{\chi} \rightarrow \mathbb{C}$ is a non-zero additive function such that $\chi^{*}=\chi$, $A \circ \sigma=-A$, and $\rho: P_{\chi} \rightarrow \mathbb{C}$ is a function such that $\rho^*=-\rho$ and the condition (I) is holding. \newline
Proceeding exactly as in the previous case we show that $h=0$ on $I_{\chi}$. Then  by using similar computations to the ones of the proof of \cite[Theorem 4.3 case B (ii)]{Ajb}, we get that
\[f=\theta +\dfrac{c_1}{2}\chi A\quad\text{on}\quad S \backslash I_{\chi}, \]
where $\theta \in \mathcal{N}_{\mu}(\sigma, S)$. This occurs in part (4) with $\beta =0$ and $\alpha=\dfrac{c_1}{2}$. This completes the proof of Theorem \ref{T1}.
\end{proof}
\begin{rem}
Note that if the semigroup $S$ is not t-compatible, d'Alembert's equation \eqref{E1} may have nontrivial solutions taking arbitrary values at some points (See \cite[Example 5.3]{EB3}).
\end{rem}
\subsection*{Acknowledgment}

% ------------------------------------------------------------------------
\end{document}